OpenStudy (anonymous):

A box with a square base and no top is to be made from a square piece of carboard by cutting 7 in. squares from each corner and folding up the sides. The box is to hold 252 in3. How big a piece of cardboard is needed?

OpenStudy (anonymous):

so far I have the height as 7 and I'm stuck at a particular part of the problem

OpenStudy (anonymous):

(x-7*2)^2 (x-14)^2 7(x-14)^2=252 7(x^2-28x+196)=252 7x^2-196x-196-252=0 7x^2-196x-448=0 7(x^2-28x-64) (x^2-28x-64) and that's where I'm stuck. I don't know how to simplify the last line

OpenStudy (anonymous):